a boy is 0.25 age of his father, after 8 years the sum of both ages is 61. now what is the age of father?
Answer by fractalier(6550)

Call their ages NOW, b and f. We have
b = .25f or 4b = f
Then we have, in 8 years,
b+8 + f+8 = 61
Substituting we have
b + 8 + 4b + 8 = 61
5b + 16 = 61
5b = 45
b = 9
f = 36 years old